OpenStudy (anonymous):

computer monitor has a height of 12 inches with a diagonal measurement of 20 inches. How many inches wide is the rectangular screen?

OpenStudy (zpupster):

 do you know the Pythagorean theorem???

OpenStudy (anonymous):

no i do not please help

OpenStudy (zpupster):

 Pythagorean theorem states a^2 + b^2 = c^2

OpenStudy (zpupster):

ok I will walk you through your problem. x^2 + (12)^2 = (20)^2 expanding the known squares x^2 + 144 = 400 now subtract 144 from both sides x^2 = 256 now take the square root both sides x = 16 our width is 16

OpenStudy (anonymous):

please help me with this one Find the hypotenuse of a right triangle with legs of 9 inches and 12 inches

OpenStudy (anonymous):

is it 7?

OpenStudy (zpupster):

no use the the theorem look at the triangle I drew c is the hypontenuse (9 ) + (12) = c^2 81 + 144 =c^2 ?=c

OpenStudy (anonymous):

50,625?

OpenStudy (zpupster):

on oyur calculator add 81 + 144 when you get that answer take the square root

OpenStudy (anonymous):

15

OpenStudy (zpupster):

good job!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

thanks you have no idea how appreciative i am of your help